OpenStudy (anonymous):

(3-9i)-(5+6i) is this -2-3i ??

OpenStudy (anonymous):

yes and also

OpenStudy (anonymous):

-2-15i

OpenStudy (anonymous):

-2+3i^2

OpenStudy (anonymous):

you know complex number i^2=-1

OpenStudy (anonymous):

3-5-9i+6i=-2-3i which is also -2+3i^2

OpenStudy (radar):

@jmes You sure about that?

OpenStudy (anonymous):

so if i^2 is -1 then would it be 1 as the final answer?

OpenStudy (radar):

On what basis did you square the i tru i^2 = -1 but what is the rationale for squaring it?

OpenStudy (radar):

I see two expressions separated by a minus sign not a multiplication sign. The asker should straighten that out. Is this a subtraction problem or a multiplication problem?

OpenStudy (anonymous):

yes I am

OpenStudy (anonymous):

it based on complex number.... you should know that

OpenStudy (anonymous):

3 - 5 = -2 -9i - 6i = - 15i -2 - 15i

OpenStudy (anonymous):

that site shown complex at very top....I confirmed to you now.

OpenStudy (radar):

Assuming it is 1. a subtraction (3-9i) -(5+6i) ------- -2-15i 2. a multiplication problem: (3-9i) X -(5+6i) = (3-9i) (-5-6i)=-15-18i+45i+54i^2 =-69+27i

OpenStudy (radar):

@jmes I checked your reference link, and totally agree, but suggest you read it carefully.

OpenStudy (radar):

Yes they are complex numbers that consist of both real and imaginary components. I suggest you treat them the way they should be treated.